OpenStudy (anonymous):

Evaluate −5 − (32 +2)

OpenStudy (pawanyadav):

First solve bracket

OpenStudy (pawanyadav):

-5-(34)

OpenStudy (anonymous):

sorry, its not 22 its 2 to the second power

OpenStudy (anonymous):

3 to the second power*

OpenStudy (anonymous):

so it would be -5 -(9+2)?

OpenStudy (pawanyadav):

Also solve bracket first

OpenStudy (pawanyadav):

Yes

OpenStudy (anonymous):

okay then i got -5 - 11

OpenStudy (pawanyadav):

Next

OpenStudy (anonymous):

then -6?

OpenStudy (pawanyadav):

No both term are negative So add them with negative sign

OpenStudy (pawanyadav):

Example. -a -b=-(a+b)

OpenStudy (anonymous):

only one is negative? its negative 5 minus positive 11 right?

OpenStudy (pawanyadav):

Got it

OpenStudy (anonymous):

well when i did it your way i got positive 6. and that is not one of the answers still

OpenStudy (pawanyadav):

In my way -5-11=-(5+11)=-16

OpenStudy (anonymous):

ohhhh okay. yeah i get it now. ty :)